what is the cost per pund of a candy mixture which consist of 2 pounds of chocolate which is $4.00 per puond and 6 pounds of carob candy which cost $5.00 per pound?
what i got was 4(2)+6(5)=x for x I got 38 now am I supposed to divide by 2 because they dont give you how many pounds for the mixture.

Answer by ptaylor(1894)

OK, Let x=cost per pound
Total weight of candy(lbs)=2+6=8 lb
Total cost of candy=2*$4.00+6*$5.00=$38
So, our equation to solve is:
8*x=$38 divide each side by 8
(Note: What we have here is total number of pounds(8) time the cost per pound(x)= total cost($38).)
x=$4.75 cost per pound of the mixture
2*$4.00 + 6* $5.00=8*$4.75
$38=$38
Hope this helps---ptaylor